# SL(2N,C) Yang-Mills Theories –

### v4

@@ Abstract
We develop a four-dimensional gauge-gravity unification based on the $SL(2N,C)$ gauge theory taken in a universal Yang–Mills type setting. The accompanying tetrads are promoted to dynamical fields whose contracted invertibility condition is interpreted as a nonlinear sigma-model type length constraint. This triggers tetrad condensation and spontaneously breaks $SL(2N,C)\rightarrow SL(2,C)\times SU(N)$, lifting all noncompact directions. A special ghost-free curvature-squared Lagrangian provides a consistent quadratic sector, while an Einstein–Cartan linear curvature term is induced radiatively from fermion loops. Below the breaking scale, only a neutral tetrad associated with graviton and $SU(N)$ vector fields remain massless, whereas axial-vector and tensor fields of the entire gauge multiplet acquire heavy masses. The matter sector clearly points to a deeper elementarity of $SL(2N,C)$ spinors, which can be identified with preon constituents whose bound states form the observed quarks and leptons. Anomaly matching between preons and composites singles out $N=8$. The chain $SL(16,C)\rightarrow SL(2,C)\times SU(8)$ then naturally yields three composite quark–lepton families, while filtering out extraneous heavy states.
@@ 1 Introduction
It has long been recognized [1, 2, 3] that gravity can be cast as a gauge theory in close analogy with the gauge description of electromagnetic, weak, and strong interactions. In this language the spin connection associated with local Lorentz transformations plays the role of a gauge field, much as the Yang–Mills bosons do for internal symmetry groups. Once this parallel is taken seriously, it is natural to ask whether both gravitational and internal gauge symmetries can be embedded into a single non-compact group so that all interactions are described within a common gauge-theoretic framework that has been intensively discussed in various forms [4, 5, 6, 7, 8, 9, 10]. However, despite an attractive algebraic structure, these schemes face well-known obstacles which we consider below in the unified $SL(2N,C)$ symmetry case.
For the gravitational sector of such unification the relevant local symmetry is $SL(2,C)$, the double cover of the proper orthochronous Lorentz group, which furnishes the appropriate spinorial representation for fermions in four-dimensional spacetime. If one assumes that the fundamental spinor fields are also responsible for organizing the internal quantum numbers, it is then tempting to extend $SL(2,C)$ to a larger group $SL(2N,C)$ in which $SU(N)$ appears as an internal symmetry subgroup. In the following we refer to such $SL(2N,C)$ based constructions as hyperunified theories (HUTs), since they collect together $SL(2,C)$ gauge gravity and $SU(N)$ grand unification within a single local group. The internal $SU(N)$ piece will be called hyperflavor, in order to distinguish it from the Lorentzian $SL(2,C)$ sector which remains neutral under internal charges. The full $SL(2N,C)$ gauge multiplet includes spin-$1$ (vector and axial-vector) and spin-$2$ (bi-vector or tensor) fields, which in principle may unify Standard Model interactions and gravity—pending resolution of known challenges.
@@
The central idea in this work is that the tetrad, when treated as a genuine dynamical field, can play three simultaneous roles. First, its invertibility (orthonormality) condition can be interpreted as a nonlinear $\sigma$-model type length constraint which forces a nonzero vacuum expectation value. Second, this VEV spontaneously breaks $SL(2N,C)$ down to $SL(2,C)\times SU(N)$, such that only the $SU(N)$ vector fields remain light while the axial and tensor components become heavy. Third, in the presence of suitable fermion multiplets the same tetrad and tensor fields generate an Einstein–Cartan linear curvature term radiatively, even if such a term is absent at tree level.
@@
The paper is organized as follows. Section 2 reviews the construction of $SL(2N,C)$ gauge theories, beginning with neutral spinors and $SL(2,C)$ symmetry, and then extending to charged spinors and full $SL(2N,C)$ hyperunification. Section 3 develops the quadratic strength-tensor Lagrangians, identifying a ghost-free choice that treats all gauge submultiplets on equal footing. In Section 4 we introduce dynamical tetrads, implement the nonlinear length constraint, and analyze the spontaneous breaking $SL(2N,C)\rightarrow SL(2,C)\times SU(N)$, including the Coleman–Weinberg selection of a hyperflavor-blind vacuum and the decoupling of hyperflavored tetrad modes. Section 5 explains how an Einstein–Cartan gravity term is induced at one loop by fermion bubbles involving tetrads and tensor fields. Section 6 turns to the matter sector and shows how preon-based $SL(16,C)$ hyperunification can accommodate three composite quark-lepton families. Section 7 summarizes the main results and outlines directions for further work.
@@ 2.1 Neutral spinors – $SL(2,C)$ symmetry
We start with the $SL(2,C)$ gauge symmetry case for a neutral spinor. Consider a Dirac field $\Psi$ which transforms in the fundamental spinor representation of $SL(2,C)$, acting in the tangent (spinor) bundle at each spacetime point:
@@
$e\mathcal{L}_{G}^{(1)}=\frac{1}{2\kappa}e_{[a}^{\mu}e_{b]}^{\nu}T_{\mu\nu}^{[ab]}\text{ , \ \ }e\equiv[-\det\mathrm{Tr}(e^{\mu}e^{\nu})/4]^{-1/2}$ (13)
where $\kappa$ is a dimensionful coefficient. Identifying $\kappa$ with $8\pi/M_{Pl}^{2}$ yields the Einstein–Cartan action in Palatini formulation [3, 14]. Varying with respect to $T_{\mu}^{ab}$ expresses it algebraically in terms of the tetrad and its derivatives, and one recovers the usual Einstein action plus torsion-induced four-fermion terms. The factor $e$ ensures invariance under general four-dimensional coordinate transformations, $GL(4,R)$ as well [3]. The underlying spacetime is still globally Minkowskian; the $SL(2,C)$ gauge group acts on the tangent (spinor) space rather than on coordinates.
@@ 2.4 Tetrads linked to gravity
if it is to define a conventional spacetime metric. In other words, only a single $SU(N)$-singlet tetrad is compatible with the usual tangent bundle structure.
One can enforce this by imposing (32) as a background constraint. Alternatively, when tetrads are treated as dynamical fields, neutral tetrads may emerge via the spontaneous breaking of $SL(2N,C)$. This way, inspired by the nonlinear $\sigma$-model type length constraint
$\frac{1}{4}e_{\mu}^{aK}e_{a}^{\mu K}=1$ (33)
which, when applied to a general tetrad, drives spontaneous symmetry breaking down to $SL(2,C)\times SU(N)$. For such a vacuum configuration, tetrads written as an appropriate sum of symmetric and antisymmetric parts,
should provide an extremum of the action with $\delta_{\mu}^{a}\delta^{K0}$ acting as the effective Higgs mode, while $\eta_{\mu b}\widehat{e}^{[ab]K}$ and $\eta_{\mu b}\widetilde{e}^{\{ab\}k}$ are zero modes. Absorption of these modes induces, as we see later in Sec. 4, masses for the submultiplets of tensor and axial-vector fields, $T_{\mu}^{[ab]K}$ and $A_{\mu}^{k}$, respectively, which are associated with the broken coset generators beyond the residual $SL(2,C)\times SU(N)$ symmetry.
@@ 4 Tetrad Condensation – Lifting non-compact directions
The full $SL(2N,C)$ gauge theory contains many more degrees of freedom than are needed to describe known low-energy physics. Our strategy is to show that a dynamical tetrad, subject to the length constraint, can spontaneously break $SL(2N,C)$ and remove the unwanted non-compact directions. In this picture the tetrad not only defines the spacetime geometry but also selects the internal symmetry and filters the spin-connection multiplets that remain light. The low-energy theory then exhibits only $SL(2,C)\times SU(N)$, with $SU(N)$ vectors and a single tensor singlet, while hyperflavored axial and tensor fields are lifted to high mass.
@@ 4.1 Dynamical tetrads
where the covariant derivative is built from the full spin-connection multiplet (23). Once such a term is added, the theory automatically extends from pure $SL(2N,C)$ to its inhomogeneous version $ISL(2N,C)$, containing local hypertranslations generated by $P_{aK}$. The tetrads $e_{\mu}^{\;aK}$ now act as the corresponding gauge fields.
After rescaling the tetrad to a dimensionless field,
$e_{\mu}^{\;aK}=M_{\!}\,E_{\mu}^{\;aK}$ (47)
where $M$ is a characteristic mass scale, the nonlinear constraint becomes
$\frac{1}{4}\,e_{\mu}^{\;aK}e^{\mu K}{}_{a}=M_{\!}^{2}$ (48)
analogous to a fixed-length condition in a nonlinear $\sigma$ model. A minimal solution consistent with this constraint is
$\overline{e}_{\mu}^{\;}=M_{\!}\,\delta_{\mu}^{\;a}\,\gamma_{a}\,\lambda^{0},\qquad\lambda^{0}=\sqrt{2/N}\mathbf{1}_{N},$ (49)
@@
where $R(e)$ is the Ricci scalar built from the tetrad and $R^{2}$ denotes the curvature-squared terms. The equations of motion for the heavy spin connection enforce metricity, $D_{\mu}e_{\nu}{}^{a}=0$, up to corrections of order $p^{2}/M^{2}$, so variation with respect to $e_{\mu}{}^{a}$ reproduces Einstein’s equation plus Planck-suppressed higher-curvature and torsion contributions.
@@ 4.2 Hyperflavor-blind vacuum
The length constraint may be imposed via a Lagrange multiplier,
[…]
which fixes the norm but leaves the internal orientation undetermined. A convenient parametrization of the vacuum is

### v5

# SL(2N,C) Yang-Mills Theories :

@@ Abstract
We develop a four-dimensional gauge-gravity unification based on the $SL(2N,C)$ gauge theory taken in a universal Yang–Mills type setting. The accompanying tetrads are promoted to dynamical fields whose length, when projected onto the background Minkowskian spacetime, is restricted by a nonlinear sigma-model type constraint. This triggers tetrad condensation and spontaneous symmetry breaking, $SL(2N,C)\rightarrow SL(2,C)\times SU(N)$, lifting all noncompact directions. A special ghost-free curvature-squared Lagrangian provides a consistent quadratic sector, while an Einstein–Cartan linear curvature term is induced radiatively from fermion loops. Below the breaking scale, only a neutral tetrad associated with graviton and $SU(N)$ vector fields remain massless, whereas axial-vector and tensor fields of the entire gauge multiplet acquire heavy masses. The matter sector clearly points to a deeper elementarity of $SL(2N,C)$ spinors, which can be identified with preon constituents whose bound states form the observed quarks and leptons. Anomaly matching between preons and composites singles out $N=8$. The chain $SL(16,C)\rightarrow SL(2,C)\times SU(8)$ then naturally yields three composite quark–lepton families, while filtering out extraneous heavy states.
@@ 1 Introduction
It has long been recognized [1, 2, 3] that gravity can be cast as a gauge theory in close analogy with the gauge description of electromagnetic, weak, and strong interactions. In this language the spin connection associated with local Lorentz transformations plays the role of a gauge field, much as the Yang–Mills bosons do for internal symmetry groups. Once this parallel is taken seriously, it is natural to ask whether both gravitational and internal gauge symmetries can be embedded into a single non-compact group so that all interactions are described within a common gauge-theoretic framework that has been intensively discussed in various forms [4, 5, 6, 7, 8, 9]. However, despite an attractive algebraic structure, these schemes face well-known obstacles which we consider below in the unified $SL(2N,C)$ symmetry case.
For the gravitational sector of such unification the relevant local symmetry is $SL(2,C)$, the double cover of the proper orthochronous Lorentz group, which furnishes the appropriate spinorial representation for fermions in four-dimensional spacetime. If one assumes that the fundamental spinor fields are also responsible for organizing the internal quantum numbers, it is then tempting to extend $SL(2,C)$ to a larger group $SL(2N,C)$ in which $SU(N)$ appears as an internal symmetry subgroup. In the following we refer to such $SL(2N,C)$ based constructions as hyperunified theories (HUTs), since they collect together $SL(2,C)$ gauge gravity and $SU(N)$ grand unification within a single local group. The internal $SU(N)$ piece will be called hyperflavor, in order to distinguish it from the Lorentzian $SL(2,C)$ sector which remains neutral under internal charges. The full $SL(2N,C)$ gauge multiplet includes spin-$1$ (vector and axial-vector) and spin-$2$ (bi-vector or tensor) fields, which in principle may unify Standard Model interactions and gravity—pending resolution of known challenges [9].
@@
The central idea in this work is that the tetrad, when treated as a genuine dynamical field, can play four simultaneous roles. First, it is, as usual, the soldering tool which connects the rigid spinorial tangent space with the flexible spacetime manifold and thereby determines the true metric structure of spacetime. Second, its background-projected length is restricted by a nonlinear $\sigma$-model type constraint [11], which enforces the tetrad condensation. Third, in the enlarged $SL(2N,C)$ framework, this spontaneously breaks $SL(2N,C)$ down to $SL(2,C)\times SU(N)$, with only the $SU(N)$ vector fields of the original gauge multiplet remaining light, while the axial and tensor submultiplets become heavy. Fourth, in the presence of suitable fermion multiplets, the same tetrad and tensor fields can radiatively generate an Einstein–Cartan linear curvature term, even if such a term is absent at tree level.
@@
The paper is organized as follows. Section 2 reviews the construction of $SL(2N,C)$ gauge theories, beginning with neutral spinors and $SL(2,C)$ symmetry, and then extending to charged spinors and full $SL(2N,C)$ hyperunification. Section 3 develops the quadratic strength-tensor Lagrangians, identifying a ghost-free choice that treats all gauge submultiplets on equal footing. In Section 4 we introduce dynamical tetrads, implement their nonlinear length constraint, and analyze the spontaneous breaking $SL(2N,C)\rightarrow SL(2,C)\times SU(N)$, including the Coleman–Weinberg selection of a hyperflavor-blind vacuum and the decoupling of hyperflavored tetrad modes. Section 5 explains how an Einstein–Cartan gravity term is induced at one loop by fermion bubbles involving tetrads and tensor fields. Section 6 turns to the matter sector and shows how preon-based $SL(16,C)$ hyperunification can accommodate three composite quark-lepton families. Section 7 summarizes the main results and outlines directions for further work.
@@ 2.1 Neutral spinors – $SL(2,C)$ symmetry
We start with the $SL(2,C)$ gauge symmetry case for a neutral spinor. Consider a Dirac field $\Psi$ which transforms in the fundamental spinor representation of $SL(2,C)$, acting in the tangent (spinor) bundle at each spacetime point [3],
@@
$e\mathcal{L}_{G}^{(1)}=\frac{1}{2\kappa}ee_{[a}^{\mu}e_{b]}^{\nu}T_{\mu\nu}^{[ab]}\text{ , \ \ }e\equiv[-\det\mathrm{Tr}(e^{\mu}e^{\nu})/4]^{-1/2}$ (13)
where $\kappa$ is a dimensionful coefficient. Identifying $\kappa$ with $8\pi/M_{Pl}^{2}$ yields the Einstein–Cartan action in Palatini formulation [3, 14]. Varying with respect to $T_{\mu}^{ab}$ expresses it algebraically in terms of the tetrad and its derivatives, and one recovers the usual Einstein action plus torsion-induced four-fermion terms. The factor $e$ ensures invariance under general four-dimensional diffeomorphisms as well, while the $SL(2,C)$ gauge group solely acts in the tangent spinorial space [3].
@@ 2.4 Tetrads linked to gravity
if it is to define a conventional spacetime metric. In other words, only a single $SU(N)$-singlet tetrad is compatible with the usual tangent bundle structure, thus remaining the metric determination (7) as in the pure $SL(2,C)$ gravity.
One can enforce this by imposing (32) as a background constraint. Alternatively, when tetrads are treated as dynamical fields, neutral tetrads may emerge via the spontaneous breaking of $SL(2N,C)$, as we see later in Sec. 4.
@@ 4 Tetrad Condensation – Lifting non-compact directions
The full $SL(2N,C)$ gauge theory contains many more degrees of freedom than are needed to describe known low-energy physics. Our strategy is to show that a dynamical tetrad, subject to the length constraint, can spontaneously break $SL(2N,C)$ and remove the unwanted non-compact directions. In this picture the tetrad not only defines the spacetime geometry but also selects the internal symmetry and filters the spin-connection multiplets that remain light. The low-energy theory then exhibits only $SL(2,C)\times SU(N)$ symmetry with massless $SU(N)$ vectors, while hyperflavored axial and tensor fields are lifted to high masses.
@@ 4.1 Dynamical tetrads
where the covariant derivative is built from the full spin-connection multiplet (23), and the whole theory indeed still admits diffeomorphism gauge invariance.
After rescaling the tetrads to a dynamical dimensionful fields,
$e_{\mu}^{\;}=M_{\!}\,E_{\mu}^{\;},\text{ }e_{\mu}^{\;aK}=M_{\!}\,E_{\mu}^{\;aK},\text{ }K=0,...,N^{2}-1$ (45)
where $M$ is a characteristic mass scale (and $E_{\mu}^{\;}$ is the ”standard” dimensionless tetrad), one comes to the modified invertiblity conditions
$e_{\mu}^{aK}e_{a}^{\nu K}=M^{2}\delta_{\mu}^{\nu}\text{, \ }e_{\mu}^{aK}e_{b}^{\mu K}=M^{2}\delta_{b}^{a}$ (46)
and metric tensor
$g_{\mu\nu}=\frac{1}{8M^{2}}\mathrm{Tr}(e_{\mu}e_{\nu})=\frac{e_{\mu}^{aK}e_{\nu}^{bK}\eta_{ab}}{M^{2}}$ (47)
where we have used the trace conditions for $\gamma$ and $\lambda$ matrices.
Remarkably, the $SL(2N,C)$ invariance still allows to put one more constraint on tetrads fixing their length in the background Minkowskian spacetime,
[…]
or, if written in components,
$\frac{1}{4}\,e_{\mu}^{\;aK}e_{\nu}^{bK}{}\eta^{\mu\nu}\eta_{ab}=M_{\!}^{2}$ (49)
which is, in fact, analogous to such a condition in a nonlinear $\sigma$ model [11]. For the metric $g_{\mu\nu}$ this reads as
$\frac{1}{4}g_{\mu\nu}\eta^{\mu\nu}=1$ (50)
which can then be checked as a gauge condition for the metric sector. On 4D Minkowski spacetime, in the smooth sector continuously connected to $g=\eta$, it is, in fact, a partial diffeomorphism gauge choice.
At the same time, the constraint (48) drives spontaneous symmetry breaking of $SL(2N,C)$ down to $SL(2,C)\times SU(N)$. For such a vacuum configuration, tetrads written as an appropriate sum of symmetric and antisymmetric parts,
should provide an extremum of the action with $\delta_{\mu}^{a}\delta^{K0}$ acting as the effective Higgs mode, while $\eta_{\mu b}\widehat{e}^{[ab]K}$ and $\eta_{\mu b}\widetilde{e}^{\{ab\}k}$ are zero modes. Absorption of these modes induces masses for the submultiplets, which are associated with the broken coset generators beyond the residual $SL(2,C)\times SU(N)$ symmetry.
Actually, the nonzero tetrad VEV matrix has a form
$\overline{e}_{\mu}^{\;}=M_{\!}\,\delta_{\mu}^{\;a}\,\gamma_{a}\,\lambda^{0},\text{ \ \ \ }\lambda^{0}=\sqrt{2/N}\mathbf{1}_{N},$ (52)
@@
where $R(e)$ is the Ricci scalar built from the tetrad and $R^{2}$ denotes the curvature-squared terms. The equations of motion for the heavy spin connection enforce metricity, $D_{\mu}e{}=0$, up to corrections of order $p^{2}/M^{2}$, so variation with respect to $e$ reproduces Einstein’s equation plus Planck-suppressed higher-curvature and torsion contributions.
@@ 4.2 Hyperflavor-blind vacuum
The length constraint (48) may be imposed via a Lagrange multiplier, which fixes the norm of teetrad but leaves its internal orientation undetermined. A convenient parametrization of the vacuum is